OpenStudy (anonymous):

Carrie is driving home to visit her parents. Which of the following statements does not describe a proportional relationship between the time Carrie spends driving and the distance she travels? A.Carrie drives 120 miles after two hours, 240 miles after 4 hours, and 360 miles after 6 hours. B.Carrie drives 50 miles after one hour, 100 miles after 2 hours, and 150 miles after 3 hours. C.Carrie drives 70 miles after one hour, 140 miles after 2 hours, and 210 miles after 4 hours. D.Carrie drives 100 miles after two hours, 200 miles after 4 hours, and 300 miles after 6 hours.

OpenStudy (anonymous):

Let's turn these into fractions. Simplify each of the following fractions; if one is different from the others in that option, then that option does not show a proportional relationship. \[A \frac{120}{2}, \frac{240}{4}, \frac{360}{6}\] \[B \frac{50}{1}, \frac{100}{2}, \frac{150}{3}\] \[C \frac{70}{1}, \frac{140}{2}, \frac{210}{4}\] \[D \frac{100}{2}, \frac{200}{4}, \frac{300}{6}\]